OpenStudy (anonymous):

Find the third iteration of f(m) = 0.5m +1 when m0 = 10 + 8i. a. 3 + i b.2.5 + 0.5(i) c.4 + 2i

OpenStudy (loser66):

first iteration of f (m) is f (0) = 0.5 * (10+8i) +1 = 6 +4i second iteration of f (m) is f (6+4i) = 0.5* (6+4i) +1 = 4+2i third time around is f9(4+2i) = 0.5*(4+2i) +1= 3 +i

OpenStudy (loser66):

so, a is my choice

OpenStudy (anonymous):

THANKS SOOO MUCH!!!!

OpenStudy (loser66):

np